# Suzhou 215006, China

### v1

@@ 3.4 Multi-Scale Merging
In a multi-scale setting, we parallelize the Token Projection Blocks and Contextual Sampling Blocks with varying token lengths and sampling rates to assure temporal patterns across different scales are properly captured. The outputs of multiple branches are fused through a convolution-based merging approach. We define $\mathbf{S}=\{(TL_{i},SR_{j})\}$ as the ensemble of parameters for all such branches containing a Token Projection Block and a Contextual Sampling Block (values are set separately for simplicity). Multiple branches corresponding to the pairs defined in $\mathbf{S}$ can be computed in parallel, generating coarse-grained representations at different scales.
Since each branch preserves a granularity of information at a different scale, the coarse-grained representations obtained from different branches are decoded independently. Specifically, separate predictors are employed to decode the coarsened representation and formulate the future prediction for each branch
@@
where $\operatorname{Predictor}(\cdot)$ is a 2-layer perceptron. $\mathbf{Y_{m}^{i,j}}\in\mathbb{R}^{(I+O)\times N}$ serves as the future prediction of a given branch. We utilize 2D convolution to blend the branch-wise predictions $\mathbf{Y_{m}^{i,j}}$, which are truncated to the desired output length. With the given parameters $\mathbf{S}$, the final prediction is computed as
@@ 4.1.2 Baseline Models and Setup
We choose representative models from three categories as our baselines, including the CNN-based model TimesNet [20], the Transformer-based models PatchTST [12], FEDformer [25], Autoformer [21], and the Linear- and MLP-based models DLinear [22] and LightTS [23]. Notably, PatchTST utilizes tokens (patches) as the transferred input in a similar manner as with our Token Projection Block.
@@
4.1.3 Main Results
For multivariate long-term time series forecasting, our model outperforms or on par with the baselines on all benchmarks, as displayed in Table 2. In comparison with the current SOTA model Transformer-based PatchTST, our model with the simple architecture demonstrates superior performance with a 4.1% reduction in MSE and a 3.3% reduction in MAE. On large datasets such as Traffic and Electricity, the proposed model consistently outperforms PatchTST in all settings. Compared with the previous best CNN-based model TimesNet, our model achieves a 6.6% reduction in MSE and a 3.3% reduction in MAE. Notably on the largest Traffic dataset, our model’s MSE is 18.2% lower. In contrast to the top-performing linear model DLinear, our model surpasses it significantly with a 14.4% decrease in MSE and an 11.6% decrease in MAE.
@@ 4.1.3 Main Results
4.2 Ablation Study
  • CP-Net: represents the standard model we propose.
  • w/o TP: represents removing coarsening in the TP block.
  • w/o CS: represents removing coarsening in the CS block.
@@ 4.2 Ablation Study
We carry out the ablation study on the three largest datasets, Electricity, Traffic and Weather, and the results are shown in Table 3. Overall, the two coarsening modules together significantly improve the performance compared with the raw MLP model, with respective improvements of 10.3%, 17.6%, and 6.5% in MSE on the three datasets. Using either one independently can enhance the performance of the MLP model, which demonstrate the importance of enhancing MLP model with the effective extraction of short-term patterns. Yet, it is noticeable that while incorporating the coarsening in the TP contributes to the performance, the coarsening module in the CS block exhibits a more pronounced effect. The reason may be attributed to the fact that this module not only captures periodic patterns through dilated and equispaced convolutions but also retains information by eliminating noise through the down-sampling process.
@@ 4.3 Model Analyses
Previous work [22, 12] shows that some Transformer-based models such as Autoformer and FEDformer degrade in the accuracy of prediction as the look-back window expands. With a longer look-back window, more useful information is exposed, such as long-term dependencies which cannot be captured within shorter ones. Therefore, a good time series prediction model should be able to make more accurate prediction as the look-back window expands. To further examine whether our model has this desirable ability, we train our model with different look-back windows $I\in\{48,96,192,336,720\}$ and compare its performance with other state-of-the-art models. The results are shown in Fig. 3. Our model performs excellently across varying look-back windows, with its performance consistently improving as the look-back window expands. This further validates the model’s capability to effectively utilize both short- and long-term information. Moreover, we can observe that TimesNet fails to achieve better performance when look-back window expands, which demonstrate the idea 2D convolution may still lack the essence of modeling long-term dependencies. At the same time, both PatchTST and DLinear exhibit subpar performance when the look-back window is small. This implies that attention and purely linear mapping may struggle to capture short-term patterns with a good accuracy.
@@
Beyond excelling performance compared with baseline models, we further demonstrate the high efficiency of the proposed model attributed to its MLP-based architecture. In this regard, we compare CP-Net with the SOTA model PatchTST and focus on the training and inference time. For fair comparisons, we conduct experiments using the data loader in Autoformer [21]. We utilize a batch size of 8 for the Electricity dataset, resulting in batches of size 8 × 321 × I (with I being the varying width of the look-back window). Specifically, the Electricity dataset consists of 321 individual time series. We report the inference time per batch and the training time for one epoch for CP-Net and PatchTST respectively as the look-back window (I) varies from 96 to 2880, as shown in Fig. 4. Our model is significantly superior to PatchTST regarding both training and inference speeds. Besides, it has smaller increase in the training and inference time with the look-back window. This is in contrast with PatchTST which exhibits a greater sensitivity to the width of the look-back window (it runs out of memory when $I\geq 2880$). It can thus be concluded that with the architectural simplicity (based purely on MLPs and convolution layers of several types), our model achieves a superior or comparable accuracy with improved computational and memory efficiency compared to models based on the attention mechanism. All experiments for this runtime comparison were conducted using a single NVIDIA RTX3090Ti GPU on the same machine.
@@ 5 Conclusion
We present CP-Net, a novel long-term multivariate time series forecasting model that effectively enhances the point-wise global projection of the MLP layer by incorporating characteristic short-term dependencies. At the core of CP-Net lies a multi-scale coarsening strategy composed of Token Projection Blocks and Contextual Sampling Blocks that integrate functions prior and posterior to the MLP. Through carefully design, CP-Net restores semantic patterns and long-range contextual patterns, which is to a large extent deficient with a sole MLP layer. In comparison with typical convolution-based models lacking the ability to model long-term dependencies and Transformer-based models with a higher computational complexity, our model is able to comprehend both local and global correlations in the time series while maintains a linear computational complexity. With this architectural simplicity, the extensive experimental results show that CP-Net outperforms or is on par with the SOTA models across nearly all the empirical datasets, which demonstrates the effectiveness of the proposed coarsening strategy for enhancing MLPs in modeling time series data.

### v2

@@ 3.4 Multi-Scale Merging
With a multi-scale setting, we integrate temporal patterns at all granularity levels. We parallelize the Token Projection Blocks and Contextual Sampling Blocks in multiple branches corresponding to diverse token lengths and sampling rates. The outputs are fused through a convolution-based merging approach. We define $\mathbf{S}=\{(TL_{i},SR_{j})\}$ as the ensemble of parameters for all such branches containing a Token Projection Block and a Contextual Sampling Block (values are set separately for simplicity). Multiple branches corresponding to the pairs defined in $\mathbf{S}$ can be computed in parallel, generating temporal patterns at different scales. Since information granules are preserved in different branches, they are decoded independently. Specifically, separate predictors are employed to decode the temporal patterns in each branch
@@
where $\operatorname{Predictor}(\cdot)$ is a 2-layer perceptron. $\mathbf{X_{m}^{i,j}}\in\mathbb{R}^{M\times N},M=(I+O)/SR_{j}$ is the output of a given branch of the Contextual Sampling Block. $\mathbf{Y_{m}^{i,j}}\in\mathbb{R}^{(I+O)\times N}$ represents a corresponding future prediction. We utilize 2D convolution to blend the branch-wise predictions $\mathbf{Y_{m}^{i,j}}$ with proper weights, which are truncated to the desired output length $O$ and compared with the ground truth in a supervised training. With the given parameters $\mathbf{S}$, the final prediction is computed as
@@ 4.1.2 Baseline Models and Setup
We choose representative models from three categories as our baselines, including the CNN-based model TimesNet [12], the Transformer-based models PatchTST [6], FEDformer [18], Autoformer [13], and the Linear- and MLP-based models DLinear [15] and LightTS [16].
4.1.3 Main Results
For multivariate long-term time series forecasting, our model outperforms or on par with the baselines on all benchmarks, as displayed in Table 2. In comparison with the current SOTA model Transformer-based PatchTST, our model with the simple architecture demonstrates superior performance with a 4.1% reduction in MSE and a 3.3% reduction in MAE. On large datasets such as Traffic and Electricity, the proposed model consistently outperforms PatchTST in all settings. Compared with the previous best CNN-based model TimesNet, our model achieves a 6.6% reduction in MSE and a 3.3% reduction in MAE. Notably on the largest Traffic dataset, our model’s MSE is 18.2% lower. In contrast to the top-performing linear model DLinear, our model surpasses it significantly with a 14.4% decrease in MSE and an 11.6% decrease in MAE. This clearly demonstrates that our coarsening strategy has effectively alleviated the drawbacks of point-wise projections and thus significantly outperforms the model in the same category.
@@ 4.1.3 Main Results
4.2 Ablation Study
  • $\bullet$
    CP-Net: represents the standard model we propose.
  • $\bullet$
    w/o TP: represents removing coarsening in the TP block.
  • $\bullet$
    w/o CS: represents removing coarsening in the CS block.
  • $\bullet$
@@ 4.2 Ablation Study
We carry out the ablation study on the three largest datasets, Electricity, Traffic and Weather, and the results are shown in Table 3. Overall, the two coarsening modules together significantly improve the performance compared with the raw MLP model, with respective improvements of 10.3%, 17.6%, and 6.5% in MSE on the three datasets. Using either one independently can enhance the performance of the MLP model, which demonstrate the importance of enhancing MLP model with the effective extraction of short-term patterns. Yet, it is noticeable that while incorporating the coarsening in the TP contributes to the performance, the coarsening module in the CS block exhibits a more pronounced effect. The reason may be attributed to the fact that this module not only captures clear patterns through dilated and equispaced convolutions but also form an adequate information bottleneck that eliminates noises through the down-sampling process.
@@ 4.3 Model Analyses
Previous work [15, 6] shows that some Transformer-based models such as Autoformer and FEDformer degrade in the accuracy of prediction as the look-back window expands. With a longer look-back window, more useful information is exposed, such as long-term dependencies which cannot be captured within shorter ones. Therefore, a good time series prediction model should be able to make more accurate prediction as the look-back window expands. To further examine whether our model has this desirable ability, we train our model with different look-back windows $I\in\{48,96,192,336,720\}$ and compare its performance with other state-of-the-art models. The results are shown in Fig. 3. Our model performs excellently across varying look-back windows, with its performance consistently improving as the look-back window expands. Moreover, we can observe that TimesNet fails to achieve better performance when look-back window expands, which demonstrate the idea 2D convolution may still lack the essence of modeling long-term dependencies.
@@
Beyond excelling performance compared with baseline models, we further demonstrate the high efficiency of the proposed model attributed to its MLP-based architecture. In this regard, we compare CP-Net with the SOTA model PatchTST and focus on the training and inference time. For fair comparisons, we conduct experiments using the data loader in Autoformer [13]. We utilize a batch size of 8 for the Electricity dataset with 321 individual time series, resulting in batches of size 8 × 321 × I (with I being the varying width of the look-back window). We report the inference time per batch and the training time for one epoch for CP-Net and PatchTST respectively as the look-back window (I) varies from 96 to 2880, as shown in Fig. 4. Our model is significantly superior to PatchTST regarding both training and inference speeds with less fluctuation. Besides, PatchTST runs out of memory when $I\geq 2880$. It can thus be concluded that with the architectural simplicity, our model achieves a superior or comparable accuracy with improved computational and memory efficiency compared to models based on the attention mechanism. All experiments for this runtime comparison were conducted using a single NVIDIA RTX3090Ti GPU on the same machine.
@@ 5 Conclusion
We present the CP-Net, a novel time series forecasting model featured by an effective boosting strategy for MLPs. It employs a coarsening scheme that addresses two critical problems of the point-wise projection of the MLP layer, known as the deficient contextual dependencies and inadequate information bottleneck. At the core of the CP-Net lies a multi-scale coarsening strategy composed of Token Projection Blocks and Contextual Sampling Blocks that integrate functions prior and posterior to the MLP. By forming diverse information granules, the CP-Net exhibits two key advantages as being able to grasp essential patterns in the time series and meanwhile filtering out unwanted volatile information, both falling short with a sole MLP layer. In comparison with typical convolution-based models lacking the long-term modeling capacity and Transformer-based models with a higher computational complexity, our model is able to comprehend both local and global temporal correlations while maintains a linear computational complexity. With its architectural simplicity, our extensive experimental results show that CP-Net outperforms or is on par with the SOTA models across nearly all adopted empirical datasets, which demonstrates the effectiveness of the proposed coarsening strategy in alleviating the intrinsic problems of MLPs and enhancing their modeling capability.